# Mean-Force Hamiltonians from Influence Functionals

### v1

@@ V Numerical Evaluation
Finally, we implement the stochastic formulation of the quenched density (Eq. 22). The Gaussian noise $\xi(\tau)$ is generated by convolution of white noise with the covariance matrix of the discretised kernel $K(\tau)$, a standard technique for sampling coloured noise (see e.g. [39]). In the present model, the variance of the integrated field $X=\int_{0}^{\beta}d\tau\,\xi(\tau)$ is predicted to be $2\beta\lambda_{\mathrm{disc}}$. Fig. 1 confirms this scaling, plotting the sample variance extracted from the stochastic simulations against $\beta$ (panel a) and $g^{2}$ (panel b).
@@
Fig. 3 demonstrates robustness in the strong-coupling regime, where the interaction strength rivals the system energy scales. To verify the operator structure of the HMF, we fit the numerically extracted $H_{\mathrm{MF}}$ to the ansatz $c_{0}\mathbb{I}+a_{H_{S}}H_{S}+a_{P}P$. As shown in panel (b), the coefficient $a_{P}$ tracks the theoretical prediction $\lambda$ perfectly, while $a_{H_{S}}\approx 1$ confirms that the bare system Hamiltonian is not renormalized.
Figure 3: Strong-coupling benchmark. (a) Population $p_{4}$ vs $g^{2}$ at fixed $\beta$, showing excellent agreement over the full range. (b) Coefficients of the HMF in the system basis vs $g^{2}$. The projector component $P=|4\rangle\langle 4|$ scales as $\lambda$ (orange/blue), while $H_{S}$ remains unrenormalised (dotted), confirming $H_{\mathrm{MF}}=H_{S}-\lambda f^{2}$.
@@ VI Discussion and outlook
The exact result derived here is hardly novel, and in isolation can hardly justify the formal machinery used to derive it. In this regard, our treatment is incomplete. The power of the quenched density lies in its modularity, which naturally accommodates both more general environments, and system bath-couplings (where $H_{S},f]\neq 0$). In the latter case, the challenge shifts from the statistical generation of noise to the Lie-algebraic properties of the time-ordered effective propagator. The framework thus clearly delineates the statistical complexity of the bath from the dynamical complexity of the system response. In forthcoming work, we shall address these generalisations. Even in this simplified case however, we see the shape of the answer to the question with which we began. What governs equilibrium? Algebra.

### v2

@@ V Numerical Evaluation
Finally, we implement the stochastic formulation of the quenched density given by Eq. 22). The Gaussian noise $\xi(\tau)$ is generated by convolution of white noise with the covariance matrix of the discretised kernel $K(\tau)$, a standard technique for sampling coloured noise (see e.g. [52]). In the present model, the variance of the integrated field $X=\int_{0}^{\beta}d\tau\,\xi(\tau)$ is predicted to be $2\beta\lambda_{\mathrm{disc}}$. Fig. 1 confirms this scaling, plotting the sample variance extracted from the stochastic simulations against $\beta$ (panel a) and $g^{2}$ (panel b).
@@
Fig. 3 demonstrates robustness in the strong-coupling regime, where the interaction strength rivals the system energy scales. To verify the operator structure of the HMF, we fit the numerically extracted $H_{\mathrm{MF}}$ to the ansatz $c_{0}\mathbb{I}+a_{H_{S}}H_{S}+a_{P}P$. As shown in panel (b), the coefficient $a_{P}$ tracks the theoretical prediction $\lambda$ perfectly, while $a_{H_{S}}\approx 1$ confirms that the bare system Hamiltonian is not renormalised.
Figure 3: Strong-coupling benchmark. (a) Population $p_{4}$ vs $g^{2}$ at fixed $\beta$, showing excellent agreement over the full range. (b) Coefficients in the system basis vs $g^{2}$. The projector component $P_{4}=|4\rangle\langle 4|$ scales as $\lambda$, while off-diagonal coherences are exponentially suppressed. This strong-coupling projector dominance is consistent with ultrastrong-coupling analyses of the mean-force Gibbs state [18]. The bare system Hamiltonian $H_{S}$ remains unrenormalised (dotted), confirming $H_{\mathrm{MF}}=H_{S}-\lambda f^{2}$.
@@ VI Discussion and outlook
The exact result derived here is not novel, and in isolation can hardly justify the formal machinery used to derive it. In this regard, our treatment is incomplete. The power of the quenched density lies in its modularity, which naturally accommodates both more general environments, and system bath-couplings. In the latter case, the challenge shifts from the statistical generation of noise to the Lie-algebraic properties of the time-ordered effective propagator. The present framework handles this by clearly delineating the statistical complexity of the bath from the dynamical complexity of the system response, meaning it is possible - though nontrivial - to incorporate both non-commuting couplings and anharmonic baths. In forthcoming work, we shall address these generalisations. Even in this simplified case however, we see the shape of the answer to the question with which we began. What governs equilibrium? Algebra.